# Random regular graph states are complex at almost any depth

### v1

@@ 4.5 Upper bound on the variance for grid graphs
Let $d=n^{c}$ for any choice of $10<c<$ 1. Let $G_{n,d}$ be a random $d$-regular graph on $n$ vertices and let $H$ be a grid graph having $v$ vertices for any $v$ satisfying $v^{2}\cdot d=\mathcal{O}(n)$, $v=o(d^{1/3})$, and $d=\omega(n^{0.5})$. Let $Y_{H}$ denote the number of induced copies of $H$ in $G_{n,d}$. Then,
@@
After proving Lemma 32, we extend it to the case of the sparsified grid, as defined in Section 4.4. This immediately leads to the main theorem:
@@
Let $G$ be a random $d$–regular graph on $n$ vertices, with $d=n^{c}$ where $10.5<c<$ 1. Then, with probability $1-o(1)$, it contains a square grid graph on $v$ vertices, for any $v=o(n^{k}$), with $k=\mathsf{min}\left\{\frac{1-c}{2},\frac{c}{3}\right\}$, as an induced subgraph.
@@
From the calculations in Corollary 30 and Lemma 32, by applying Chebyshev’s inequality,
@@
Hence, with probability $1-o(1)$, $Y_{H}$ is non-zero, where the probability is taken over the randomness in the choice of $G_{n,d}$. This completes the proof. ∎
Now we prove Lemma 32.
Proof of Lemma 32.
@@
Let $H^{\prime}$ and $H^{\prime\prime}$ have at most one vertex in common. Then, by Lemma 28,
@@
The second line follows from Equation 171. The third line follows from Lemma 28. The fourth line follows from the fact that there are
@@
$\leq\sum_{F\subseteq H,e_{F}>0}\leavevmode\nobreak\ o\big{(}n^{2v-v_{F}}\big{)}\leavevmode\nobreak\ \mathbb{E}(J_{H^{\prime}}J_{H^{\prime\prime}})$
@@
The fourth line follows from Equation 158. Now, note that the number of subgraphs of $F$ is
@@
By the definition of the density $m(H)$, one sufficient condition for Equation 180 to hold is
@@
$\binom{n}{t}=o(n^{t}),\leavevmode\nobreak\ \leavevmode\nobreak\ \binom{n-t}{v-t}=o(n^{v-t}),\leavevmode\nobreak\ \leavevmode\nobreak\ \binom{n-v}{v-t}=o(n^{v-t}).$ (185)
Again by Lemma 28, we have that
@@
In Equation 186, we subtract $\binom{t}{2}$ edges from the exponent of the term in the extreme right because we are counting them twice — once in the complement set of $H^{\prime}$ and again in the complement set of $H^{\prime\prime}$. Hence,
@@
Let $G_{n,d}$ be a random $d$-regular graph on $n$ vertices and let $H$ be a sparsified square grid graph having $v$ vertices for any $v$ satisfying $v^{2}\cdot d=\mathcal{O}(n)$, $v=o(d^{1/3})$, and $d=n^{c}$ for any constant $10<c<$ 1. Let $Y_{H}$ denote the number of induced copies of $H$ in $G_{n,d}$. Then,
@@
The proof is the same as that of Lemma 32 with the observation that in the second case, if $H$ is the sparsified square grid graph, the local density is bounded as
@@ 5.1 Lower bound on geometric entanglement
In this section we will prove LABEL:thm:geom_ent_lower_stab and LABEL:cor:geom_ent_lower_graph. We begin by extending a nice trick from compressed sensing (see e.g. [71, Lemma 4.4.1]) on approximating extremal singular values of matrices through epsilon nets (we found essentially this argument in [68] but it is probably folklore in the tensor community).
@@
Consider an $n$ qubit state |$\ket{\psi}$ and let $A_{\epsilon/n}$ be an $(\ln(3/2)/n)$-net of the set of single qubit states. We then have that
$\max_{\ket{\alpha}\in\mathrm{PROD}}|\langle\alpha|\psi\rangle|^{2}\leq 2\max_{\ket{\beta}\in A_{\ln(3/2)/n}^{\otimes n}}|\langle\beta|\psi\rangle|^{2}.$ (193)
@@
Consider a state |$\ket{\alpha^{*}}$ such that $\max_{\ket{\alpha}\in\mathrm{PROD}}|\langle\alpha|\psi\rangle|^{2}=|\langle\alpha^{*}|\psi\rangle|^{2}$. Writing |$\ket{\alpha^{*}}=\bigotimes_{i=1}^{n}\ket{\alpha^{*}_{i}}$ we choose for each |$\ket{\alpha^{*}_{i}}$ a state |$\ket{\beta_{i}}\in A_{\ln(3/2)/n}$ such that |$\norm{\ket{\alpha^{*}_{i}}-\ket{\beta_{i}}}\leq\ln(3/2)/n$. We can write
$|\langle\alpha^{*}|\psi\rangle|=|\langle\bigotimes_{i=1}^{n}(\bra{\beta_{i}}+\bra{\alpha^{*}_{i}}-\bra{\beta_{i}})|\psi\rangle|\leq|\langle\beta|\psi\rangle|+\sum_{i=1}^{n}\binom{n}{i}(\ln(3/2)/n)^{i}|\langle\alpha^{*}|\psi\rangle|,$ (194)
@@
Recall that we are trying to lower-bound the geometric entanglement. This is equivalent to upper-bounding the maximization over product states. The probability that the geometric entanglement is small is equal to the probability that the maximization of the overlap is large. Now, we upper bound this probability. Combining the above argument with the union bound and the formula for average stabilizer states Equation 9, we obtain:
@@
Choose a stabilizer state |$\ket{S}$ uniformly at random. There exists a constant $c$ such that
$\mathbb{P}\bigg{[}E_{g}(\ket{S})\leq n-c\sqrt{n}\log(n)\bigg{]}\leq O(2^{-\sqrt{n}}).$ (196)
@@
$\mathbb{P}\Big{[}E_{g}(\ket{S})\leq n-\delta\Big{]}=\mathbb{P}\left[-\log\big(\max_{\alpha\in\mathrm{PROD}_{n}}|\innerproduct{\alpha}{\psi}|^{2}\big{missing})\leq n-\delta\right]\\ =\mathbb{P}\Big{[}\max_{\ket{\alpha}\in\mathrm{PROD}}|\langle\alpha|S\rangle|^{2}\geq 2^{-n+\delta}\Big{]}\leq\mathbb{P}\Big{[}\max_{\ket{\beta}\in A_{1/n}^{\otimes n}}|\langle\beta|S\rangle|^{2}\geq 2^{-n+\delta-1}\Big{]},$ (197)
@@
$\mathbb{P}\Big{[}\max_{\ket{\beta}\in A_{\ln(3/2)/n}^{\otimes n}}|\langle\beta|\psi\rangle|^{2}\leq 2^{-n+\delta-1}\Big{]}\leq 2^{n-\delta+1}\mathbb{E}_{S}\bigg{[}\sum_{\ket{\beta}\in A_{\ln(3/2)/n}^{\otimes n}}|\langle\beta|S\rangle|^{2t}\bigg{]}^{1/t}$ (198)
for some integer $t>1$ (we will specify this later) Next we use the concavity of $x^{1/t}$ and the duality formula described in Equation 9 to obtain
$2^{n-\delta-1}\mathbb{E}_{S}\bigg{[}\sum_{\ket{\beta}\in A_{\ln(3/2)/n}^{\otimes n}}|\langle\beta|S\rangle|^{2t}\bigg{]}^{1/t}\!\!\!\!\leq 2^{n-\delta-1}|A_{\ln(3/2)/n}|^{n/t}2^{-n/t}(2^{t-1}+1)^{1/t}\bigg{[}\prod_{i=0}^{t-2}\frac{(2^{i}+1)}{(2^{n}\!+\!2^{i})}\bigg{]}^{1/t}.$ (199)
@@
$2^{n-\delta+1}\mathbb{E}_{S}\bigg{[}\sum_{\ket{\beta}\in A_{\ln(3/2)/n}^{\otimes n}}|\langle\beta|S\rangle|^{2t}\bigg{]}^{1/t}\leq 5^{\sqrt{n}}2^{-\delta}2^{\log(n)\sqrt{n}}2^{\sqrt{n}}.$ (200)
@@
The above statement holds for uniformly random stabilizer states. We can lift it to a statement about graph states by noting that the probability that a uniformly random stabilizer state is a graph state (up to an irrelevant linear phase) is lower bounded by $e^{-2}$. Since conditioning on inclusion in a subset under a uniform distribution yields a uniform distribution on that subset we can say that
$\mathbb{P}\big{[}E_{g}(\ket{G})\leq n-\delta\big{]}\leq e^{2}\mathbb{P}\big{[}E_{g}(\ket{S})\leq n-\delta\big{]},$ (201)
from which LABEL:cor:geom_ent_lower_graph immediately follows.
@@ 5.2 Upper bound on geometric entanglement
$\mathbb{E}_{G}(E_{g}(\ket{G}))\leq\mathbb{E}_{G}\left[-\log(\max_{\ket{s}\in R_{n}}|\bra{G}s\rangle|^{2})\right]=n-\mathbb{E}_{G}\left[\max_{S\subseteq[n]}\left(|S|-\mathrm{Rk}\!\left(A_{G}[S]\right)\right)\right],$ (202)
@@
The first inequality is obvious from the definition of the geometric entanglement $E_{g}$. Now note that we can specify any state |$\ket{s}\in R_{n}$ by a set $S\subset[n]$ indicating the Hadamard basis part of the state and two bitstrings $x_{S}\in\{0,1\}^{|S|}$ and $y_{\bar{S}}\in\{0,1\}^{n-|S|}$ encoding the phase and computational basis states in $S,\bar{S}$ so that |$\ket{s}=2^{-|S|/2}\sum_{z\in\{0,1\}^{|S|}}(-1)^{x_{S}^{T}z}\ket{z}\ket{y_{\bar{S}}}$. To ease notation, let us assume wlog. that $S=\{1,\ldots,|S|\}$, i.e., the submatrix $A_{G}[S]$ is just the top-left corner of $A_{G}$. Then
$\max_{\ket{s}\in R_{n}}|\bra{G}s\rangle|^{2})=2^{-n-|S|}\max_{S\subset[n]}\max_{x_{S}\in\{0,1\}^{|S|}}\max_{y_{\bar{S}}\in\{0,1\}^{n-|S|}}\left|\sum_{z\in\{0,1\}^{|S|}}(-1)^{(z,y_{\bar{S}})^{T}U_{G}(z,y_{\bar{S}})+z^{T}x_{S}}\right|^{2},$ (203)
@@
The final term in this equation contributes only a global phase, while the middle one can be absorbed in the maximisation over $x_{S}$. Hence we get
$\max_{\ket{s}\in R_{n}}|\bra{G}s\rangle|^{2})=2^{-n-|S|}\max_{S\subset[n]}\max_{x_{S}\in\{0,1\}^{|S|}}\bigg{|}\sum_{z\in\{0,1\}^{|S|}}(-1)^{z^{T}{U}_{SS}z+x_{S}^{T}z}\bigg{|}^{2}.$ (205)
Next we use a theorem due to Dickson [50, Ch. 15 Thm. 2], which tells us that there exists an invertible binary matrix $P_{SS}$ such that $P_{SS}U_{SS}P^{-1}_{SS}=D_{SS}$ with
@@
$\max_{\ket{s}\in R_{n}}|\bra{G}s\rangle|^{2})=2^{-n-|S|}\max_{S\subset[n]}\bigg{(}\max_{y_{1},y_{2}\in\{0,1\}}\bigg{|}\sum_{a_{1},a_{2}\in\{0,1\}}(-1)^{a_{1}a_{2}+y_{1}a_{1}+y_{2}a_{2}}\bigg{|}^{2}\bigg{)}^{\mathrm{rank}(U_{SS})/2}.$ (208)
@@
The formula above, found in [50, Ch.15 Thm.2] [^1] Note however that this equation in [50] contains a typo, we provide the correct formula., can be both upper and lower bounded by a Gaussian. We have the following lemma, which follows from a straightforward calculation.
@@
Consider a uniformly random adjacency matrix $A$ as in 2. We have the following approximation:
@@
Consider a uniformly random $n\times n$ adjacency matrix $A$ and two sets of indices $S,I\subseteq[n]$, with $I\subseteq S$. Also consider integers $i,j$ such that $|S|-i,|I|-j$ are even. There exists a constant $\rho<1-10^{-5}$ such that
@@
because $v^{T}x=x^{T}Ax=0$ by the symmetry of $A$ and the fact that we are working over the field $\mathbb{F}_{2}$. This implies that $\mathrm{Rk}\!\left(\begin{pmatrix}0&v^{T}\\ v&A\end{pmatrix}\right)=r$ (this elegant argument is due to Sloane and MacWilliams [50, Ch. 15 Lem. 3]). The sequential adding of random $k$ columns (and rows) to a matrix $A$ of rank $r$, and considering their rank, thus induces a sequence of random variables $R_{0}=r,R_{1},\ldots,R_{k}$. From the above discussion, this sequence is a (time dependent) Markov chain with transitions $r\xrightarrow{2^{r-m}}r$, $r\xrightarrow{1-2^{r-m}}r+2$ (where $m$ takes the role of time).
@@
for $i,j\in\mathbb{N}$. This Markov chain is irreducible, but is periodic with period $2$. Hence it is natural to consider $Q=M^{2}$, which will be aperiodic, but decomposes into even and odd irreducible aperiodic subchains. We will now bound the convergence of the Markov chain $Q$ with initial state $e_{r}$ (the unit vector with $1$ on the $r$’th position). We will treat the even subchain in detail (with the odd subchain being analogous), so assume that $r$ is even. The stationary state $\pi$ of $M^{2}$ on the even subspace can be found by appropriately taking the limit of Equation 210 to $n\to\infty$:
@@
where $C\geq 1$ is some appropriate normalization. The above makes intuitive sense, because it corresponds to the situation where $|S|$ is much larger than $|I|$, and thus the rank distribution of $A[|S|]$ stops depending on $I$. One can also explicitly verify that $\pi$ is an eigenvector of $M^{2}$ with eigenvalue $1$. We will use the drift and minorization method (see Theorem 16) to bound convergence to this distribution. For this bounding method we need to provide a drift function $V:\mathbb{N}/2\to\mathbb{R}$ and a small set $C\subset\mathbb{N}/2$. We will choose $V(i)=2^{i}$ and $C=\{0,2\}$. It is tedious but straightforward to check that $P^{2}$ satisfies the conditions of Theorem 16 with parameters $\lambda=0.55,b=2,\delta=0.2,d=9,r=0.001$ (all the difficulty lies in choosing the parameters). This bounds the convergence of the Markov chain in total variation distance as
@@
[…]
@@
Choose a graph state |$\ket{G}$ on $n$ qubits uniformly at random, where we assume $n=k^{2}$ for some integer $k$. We have
$\mathbb{E}_{G}\big{(}E_{g}(\ket{G}\big{)}\leq n-\Omega(n^{1/4}/\log(n)).$ (221)
@@
We begin by upper-bounding the geometric entanglement of a graph state |$\ket{G}$ in terms of the maximal rank deficiency of the principal submatrices of the adjacency matrix of $G$. Lemma 37 tells us that
$\mathbb{E}_{G}\big{(}E_{g}(\ket{G})\big{)}\leq n-\mathbb{E}_{G}\left[\max_{S\subseteq[n]}\big{(}|S|-\mathrm{Rk}\!\left(A_{G}[S]\right)\big{)}\right].$ (222)
@@
Since $|S|=n/2$ for $S\in\mathcal{Q}$, we can lower bound the expected maximal rank deficiency (over $\mathcal{Q}$) by defining the events $E_{S}(t)=\{\mathrm{Rk}(A)\leq n/2-t\}$; recall Lemma 39. From the Markov and Bonferroni inequalities (Fact 1) we see
@@
We can give a lower bound of the first term (1) using Lemma 39:
@@
It remains to upper bound the second term (2). To this end, we first recall Lemma 40 which implies that the joint probability $\mathbb{P}\big{(}E_{S}(t)\cup E_{S^{\prime}}(t)\big{)}$ only depends on the size of the intersection $I=S\cap S^{\prime}$. The size of this intersection can take values $|I|=0,\sqrt{n},\ldots,n/2-\sqrt{n}$. Given a set $S\in\mathcal{Q}$, for each value of $w\in[0,\sqrt{n}/2-1]$ there are $\binom{\sqrt{n}}{w}\binom{\sqrt{n}}{\sqrt{n}-w}$ sets $S^{\prime}\in\mathcal{Q}$ satisfying $|S\cap S^{\prime}|=w\sqrt{n}$. Let us choose representatives $S_{0}\coloneqq[n/2]$, $S^{\prime}_{w}\coloneqq[n/2-\sqrt{n}w,n-\sqrt{n}w]$ with $I_{w}\coloneqq S_{0}\cap S^{\prime}_{w}$ of size $|I_{w}|=w\sqrt{n}$. We can then write
@@
×$\hskip 284.52756pt\times\mathbb{P}\left(\mathrm{Rk}\!\left(A[I]\right)=j\right)$ (228)
@@
×$\hskip 113.81102pt\times\mathbb{P}\left(\mathrm{Rk}\!\left(A[S_{w}^{\prime}]\right)=i^{\prime}\,|\,\mathrm{Rk}\!\left(A[I_{w}]\right)=j\right)\mathbb{P}\left(\mathrm{Rk}\!\left(A[I_{w}]\right)=j\right)$ (229)
@@
×$\hskip 50.00008pt\times\sum_{j=0}^{w\sqrt{n}/2}\bigg{(}\sum_{i=j}^{n/2-t}\mathbb{P}\big{(}\mathrm{Rk}\!\left(A[S_{0}]\right)=i\;\big{|}\;\mathrm{Rk}\!\left(A[I_{w}]\right)=j\big{)}\bigg{)}^{2}\mathbb{P}\big{(}\mathrm{Rk}\!\left(A[I_{w}]\right)=j\big{)}.$ (230)
@@
Next we use 2, which tells us the rank of $A[I_{w}]$ is likely nearly maximal. To make this precise we introduce a constant $\alpha\geq 1$ which will be precisely determined later, and split the sum over $j$ into $j<|I_{w}|-\alpha t$ and $j\geq|I_{w}|-\alpha t$. Focusing on this sum we see
@@
where we’ve used 2 to upper bound $\mathbb{P}\big{(}\mathrm{Rk}\!\left(A[I]\right)=j\big{)}$ in the second term, and have trivially bounded $\mathbb{P}\big{(}\mathrm{Rk}\!\left(A[I]\right)=j\big{)}\leq 1$ in the first. It remains to use Lemma 41 to bound the first term. We obtain

### v2

@@ 4.5 Upper bound on the variance for grid graphs
Let $d=n^{c}$ for any choice of $0<c<1$. Let $G_{n,d}$ be a random $d$-regular graph on $n$ vertices and let $H$ be a grid graph having $v$ vertices for any $v$ satisfying $v^{2}\cdot d=\mathcal{O}(n)$, $v=o(d^{1/3})$, and $d=\omega(n^{0.5})$. Let $Y_{H}$ denote the number of induced copies of $H$ in $G_{n,d}$. Then,
@@
After proving Lemma˜32, we extend it to the case of the sparsified grid, as defined in Section˜4.4. This immediately leads to the main theorem:
@@
Let $G$ be a random dd–regular graph on $n$ vertices, with $d=n^{c}$ where $0.5<c<1$. Then, with probability 1o(1)1-o(1), it contains a square grid graph on $v$ vertices, for any $v=o(n^{k}$), with $k=\mathsf{min}\left\{\frac{1-c}{2},\frac{c}{3}\right\}$, as an induced subgraph.
@@
From the calculations in Corollary˜30 and Lemma˜32, by applying Chebyshev’s inequality,
@@
Hence, with probability 1o(1)1-o(1), $Y_{H}$ is non-zero, where the probability is taken over the randomness in the choice of $G_{n,d}$. This completes the proof. ∎
Now we prove Lemma˜32.
Proof of Lemma˜32.
@@
Let $H^{\prime}$ and $H^{\prime\prime}$ have at most one vertex in common. Then, by Lemma˜28,
@@
The second line follows from Equation˜171. The third line follows from Lemma˜28. The fourth line follows from the fact that there are
@@
$\leq\sum_{F\subseteq H,e_{F}>0}\penalty 10000\ o\big(n^{2v-v_{F}}\big)\penalty 10000\ \mathbb{E}(J_{H^{\prime}}J_{H^{\prime\prime}})$
@@
The fourth line follows from Equation˜158. Now, note that the number of subgraphs of $F$ is
@@
By the definition of the density $m(H)$, one sufficient condition for Equation˜180 to hold is
@@
$\binom{n}{t}=o(n^{t}),\penalty 10000\ \penalty 10000\ \binom{n-t}{v-t}=o(n^{v-t}),\penalty 10000\ \penalty 10000\ \binom{n-v}{v-t}=o(n^{v-t}).$ (185)
Again by Lemma˜28, we have that
@@
In Equation˜186, we subtract $\binom{t}{2}$ edges from the exponent of the term in the extreme right because we are counting them twice — once in the complement set of $H^{\prime}$ and again in the complement set of $H^{\prime\prime}$. Hence,
@@
Let $G_{n,d}$ be a random $d$-regular graph on $n$ vertices and let $H$ be a sparsified square grid graph having $v$ vertices for any $v$ satisfying $v^{2}\cdot d=\mathcal{O}(n)$, $v=o(d^{1/3})$, and $d=n^{c}$ for any constant $0<c<1$. Let $Y_{H}$ denote the number of induced copies of $H$ in $G_{n,d}$. Then,
@@
The proof is the same as that of Lemma˜32 with the observation that in the second case, if $H$ is the sparsified square grid graph, the local density is bounded as
@@ 5.1 Lower bound on geometric entanglement
In this section we will prove LABEL:thm:geom_ent_lower_stab and LABEL:cor:geom_ent_lower_graph. We begin by extending a nice trick from compressed sensing (see e.g. [91, Lemma 4.4.1]) on approximating extremal singular values of matrices through epsilon nets (we found essentially this argument in [87] but it is probably folklore in the tensor community).
@@
[…]
@@
[…]
@@
Recall that we are trying to lower-bound the geometric entanglement. This is equivalent to upper-bounding the maximization over product states. The probability that the geometric entanglement is small is equal to the probability that the maximization of the overlap is large. Now, we upper bound this probability. Combining the above argument with the union bound and the formula for average stabilizer states Equation˜9, we obtain:
@@
[…]
@@
[…]
@@
[…]
for some integer $t>1$ (we will specify this later) Next we use the concavity of $x^{1/t}$ and the duality formula described in Equation˜9 to obtain
[…]
@@
[…]
@@
The above statement holds for uniformly random stabilizer states. We can lift it to a statement about graph states by noting that the probability that a uniformly random stabilizer state has full support on the computational basis states is bounded from below by a constant $C$. This is a standard fact, but we prove it here for completeness:
Corollary 37 (restatement of LABEL:cor:geom_ent_lower_graph).
[…]
Proof.
Consider the stabilizer states of full support:
[…]
with $U$ a binary upper triangular matrix and $b\in\mathbb{Z}_{4}^{n}$. We will argue that all pairs $U,b$ correspond to different quantum states. For two such states we can compute the inner product:
[…]
The matrix $U+V\mod 2$ is again upper triangular, and hence a theorem due to Dickson [67, Ch. 15 Thm. 2] tells us that there exists an invertible binary matrix $P$ such that $P(U+V)P^{-1}=D$ with
$x^{T}Dx=\sum_{i=1}^{\mathrm{rank}(U+V)/2}{x}_{2i-1}{x}_{2i}.$ (204)
Note that the rank of $U+V$ is always even, so this is sensible. Absorbing the map $P$ into the summation, we can write:
[…]
In order for this overlap to be one we must have
$(-1)^{x^{T}Dx}i^{(Px)^{T}(a+b)}=1,$ (206)
for all $x\in\{0,1\}^{n}$. This immediately implies that $i^{(Px)^{T}(a+b)}=\pm 1$ and hence that $a+b=2z$ for some $z\in\{0,1\}^{n}$. Furthermore we must have $x^{T}Dx+x^{T}(P^{T}z)=0$ for all $x\in\{0,1\}^{n}$. Setting $P^{T}z=z^{\prime}$ we can write this as
$0=\sum_{i=0}^{r/2}x_{2i}x_{2i+1}+x_{2i}z^{\prime}_{2i}+x_{2i+1}z^{\prime}_{2i+1}+\sum_{i=r/2+1}^{n/2}x_{2i}z^{\prime}_{2i}+x_{2i+1}z^{\prime}_{2i+1},$ (207)
[…]
$2^{\frac{n(n-1)}{2}+2n}\bigg[2^{n}\prod_{i=1}^{n}(2^{i}+1)\bigg]^{-1}=2^{\frac{n(n-1)}{2}+2n-n-\frac{n(n+1)}{2}}\prod_{i=1}^{n}(1+2^{-i})^{-1}\geq\prod_{i=1}^{\infty}(1+2^{-i})^{-1}=C$ (208)
[…]
where in the RHS we choose $U,b$ uniformly at random. Since conditioning on inclusion in a subset under a uniform distribution yields a uniform distribution on that subset we can say that
[…]
which is what we wanted to show. ∎
@@ 5.2 Upper bound on geometric entanglement
[…]
@@
[…]
@@
The final term in this equation contributes only a global phase, while the middle one can be absorbed in the maximization over $x_{S}$. Hence we get
[…]
Next we use a theorem due to Dickson [67, Ch. 15 Thm. 2], which tells us that there exists an invertible binary matrix $P_{SS}$ such that $P_{SS}U_{SS}P^{-1}_{SS}=D_{SS}$ with
@@
[…]
@@
The formula above, found in [67, Ch.15 Thm.2] [^4] Note however that this equation in [67] contains a typo, we provide the correct formula., can be both upper and lower bounded by a Gaussian. We have the following lemma, which follows from a straightforward calculation.
@@
Consider a uniformly random adjacency matrix $A$ as in ˜2. We have the following approximation:
@@
Consider a uniformly random $n\times n$ adjacency matrix $A$ and two sets of indices $S,I\subseteq[n]$, with $I\subseteq S$. Also consider integers $i,j$ such that |S|i,|I|j|S|-i,|I|-j are even. There exists a constant $\rho<1-10^{-5}$ such that
@@
because $v^{T}x=x^{T}Ax=0$ by the symmetry of $A$ and the fact that we are working over the field $\mathbb{F}_{2}$. This implies that $\mathrm{Rk}\!\left(\begin{pmatrix}0&v^{T}\\ v&A\end{pmatrix}\right)=r$ (this elegant argument is due to Sloane and MacWilliams [67, Ch. 15 Lem. 3]). The sequential adding of random $k$ columns (and rows) to a matrix $A$ of rank $r$, and considering their rank, thus induces a sequence of random variables $R_{0}=r,R_{1},\ldots,R_{k}$. From the above discussion, this sequence is a (time dependent) Markov chain with transitions $r\xrightarrow{2^{r-m}}r$, $r\xrightarrow{1-2^{r-m}}r+2$ (where $m$ takes the role of time).
@@
for $i,j\in\mathbb{N}$. This Markov chain is irreducible, but is periodic with period $2$. Hence it is natural to consider $Q=M^{2}$, which will be aperiodic, but decomposes into even and odd irreducible aperiodic subchains. We will now bound the convergence of the Markov chain $Q$ with initial state $e_{r}$ (the unit vector with $1$ on the $r$’th position). We will treat the even subchain in detail (with the odd subchain being analogous), so assume that $r$ is even. The stationary state $\pi$ of $M^{2}$ on the even subspace can be found by appropriately taking the limit of Equation˜219 to $n\to\infty$:
@@
where $C\geq 1$ is some appropriate normalization. The above makes intuitive sense, because it corresponds to the situation where $|S|$ is much larger than $|I|$, and thus the rank distribution of $A[|S|]$ stops depending on $I$. One can also explicitly verify that $\pi$ is an eigenvector of $M^{2}$ with eigenvalue $1$. We will use the drift and minorization method (see Theorem˜16) to bound convergence to this distribution. For this bounding method we need to provide a drift function $V:\mathbb{N}/2\to\mathbb{R}$ and a small set $C\subset\mathbb{N}/2$. We will choose $V(i)=2^{i}$ and $C=\{0,2\}$. It is tedious but straightforward to check that $P^{2}$ satisfies the conditions of Theorem˜16 with parameters $\lambda=0.55,b=2,\delta=0.2,d=9,r=0.001$ (all the difficulty lies in choosing the parameters). This bounds the convergence of the Markov chain in total variation distance as
@@
$\mathbb{P}\big(\mathrm{Rk}\!\left(A[S]\right)\leq|S|-i\;\big|\;\mathrm{Rk}\!\left(A[I]\right)=|I|-j\big)]\leq 7\rho^{|S|-|I|}2^{j}+\sum_{t=i}^{\infty}\pi_{t}=7\rho^{|S|-|I|}2^{j}+2\,2^{-\frac{i^{2}}{2}+\frac{i}{2}},$ (229)
@@
[…]
@@
[…]
@@
Since $|S|=n/2$ for $S\in\mathcal{Q}$, we can lower bound the expected maximal rank deficiency (over $\mathcal{Q}$) by defining the events $E_{S}(t)=\{\mathrm{Rk}(A)\leq n/2-t\}$; recall Lemma˜40. From the Markov and Bonferroni inequalities (Fact 1) we see
@@
We can give a lower bound of the first term (1) using Lemma˜40:
@@
It remains to upper bound the second term (2). To this end, we first recall Lemma˜41 which implies that the joint probability $\mathbb{P}\big(E_{S}(t)\cup E_{S^{\prime}}(t)\big)$ only depends on the size of the intersection $I=S\cap S^{\prime}$. The size of this intersection can take values $|I|=0,\sqrt{n},\ldots,n/2-\sqrt{n}$. Given a set $S\in\mathcal{Q}$, for each value of $w\in[0,\sqrt{n}/2-1]$ there are $\binom{\sqrt{n}}{w}\binom{\sqrt{n}}{\sqrt{n}-w}$ sets $S^{\prime}\in\mathcal{Q}$ satisfying $|S\cap S^{\prime}|=w\sqrt{n}$. Let us choose representatives $S_{0}\coloneqq[n/2]$, $S^{\prime}_{w}\coloneqq[n/2-\sqrt{n}w,n-\sqrt{n}w]$ with $I_{w}\coloneqq S_{0}\cap S^{\prime}_{w}$ of size $|I_{w}|=w\sqrt{n}$. We can then write
@@
$\hskip 284.52756pt\times\mathbb{P}\left(\mathrm{Rk}\!\left(A[I]\right)=j\right)$ (237)
@@
$\hskip 113.81102pt\times\mathbb{P}\left(\mathrm{Rk}\!\left(A[S_{w}^{\prime}]\right)=i^{\prime}\,|\,\mathrm{Rk}\!\left(A[I_{w}]\right)=j\right)\mathbb{P}\left(\mathrm{Rk}\!\left(A[I_{w}]\right)=j\right)$ (238)
@@
$\hskip 50.00008pt\times\sum_{j=0}^{w\sqrt{n}/2}\bigg(\sum_{i=j}^{n/2-t}\mathbb{P}\big(\mathrm{Rk}\!\left(A[S_{0}]\right)=i\;\big|\;\mathrm{Rk}\!\left(A[I_{w}]\right)=j\big)\bigg)^{2}\mathbb{P}\big(\mathrm{Rk}\!\left(A[I_{w}]\right)=j\big).$ (239)
@@
Next we use ˜2, which tells us the rank of $A[I_{w}]$ is likely nearly maximal. To make this precise we introduce a constant $\alpha\geq 1$ which will be precisely determined later, and split the sum over $j$ into $j<|I_{w}|-\alpha t$ and $j\geq|I_{w}|-\alpha t$. Focusing on this sum we see
@@
where we’ve used ˜2 to upper bound $\mathbb{P}\big(\mathrm{Rk}\!\left(A[I]\right)=j\big)$ in the second term, and have trivially bounded $\mathbb{P}\big(\mathrm{Rk}\!\left(A[I]\right)=j\big)\leq 1$ in the first. It remains to use Lemma˜42 to bound the first term. We obtain